### v1

# Season’s Greetings by AD

Abstract
We use Algorithmic Differentiation (AD) [1] to implement type-generic tangent and adjoint versions of
$y=\sum_{i=0}^{n-1}x_{2i}\cdot x_{2i+1}$
in C++. We run an instantiation for char-arithmetic and we print the gradient at $(101~{}77~{}114~{}114~{}32~{}121~{}109~{}88~{}115~{}97)^{T}$ to std::cout, yielding the output “Merry Xmas”.
Similar instantiations of type-generic second-order tangent and second-order adjoint versions of
$y=\frac{1}{6}\cdot\sum_{i=0}^{n-1}x^{3}_{i}$
yield “Happy 2024” at $(72~{}97~{}112~{}112~{}121~{}32~{}50~{}48~{}50~{}52)^{T}.$ Prepend a sufficiently large number of zeros to the input vector to explore the varying run times of the different derivative codes. The entire source code can be found on
https://github.com/un110076/SeasonsGreetings .
1 AD wishes Merry Xmas!
1.1 The Maths
The gradient of the function $f:{\mathbb{R}}^{N}\rightarrow{\mathbb{R}}:y=f({\bf x}),$
$y=\sum_{i=0}^{n-1}x_{2i}\cdot x_{2i+1}\;,$
where ${\bf x}=(x_{j})_{j=0}^{N-1}$ and $N=2n$ for given $n\geq 0,$ is easily found to be equal to
$f^{\prime}({\bf x})=\left(\left(\begin{pmatrix}x_{2i+1}\\ x_{2i}\end{pmatrix}\right)_{i=0}^{n-1}\right)^{T}\in{\mathbb{R}}^{1\times N}\;.$
For example, $f^{\prime}({\bf x})=(1~{}2~{}3~{}4)$ at ${\bf x}=(2~{}1~{}4~{}3)^{T}.$
A (first-order) tangent version
$f^{(1)}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\rightarrow{\mathbb{R}}:y^{(1)}=f^{(1)}({\bf x},{\bf x}^{(1)})$
of $f$ resulting from tangent AD applied to a given differentiable implementation of $f$ computes
$y^{(1)}\equiv f^{\prime}({\bf x})\cdot{\bf x}^{(1)}$
without explicit accumulation of the gradient. The latter results from letting ${\bf x}^{(1)}$ range over the Cartesian basis vectors ${\bf e}_{i}\in{\mathbb{R}}^{N},$ resulting in $N$ calls of ${f}^{(1)}$ for $i=0,\ldots,N-1.$ For example,
$y^{(1)}=f^{\prime}({\bf x})\cdot{\bf e}_{2}=(1~{}2~{}3~{}4)\cdot\begin{pmatrix}0\\ 0\\ 1\\ 0\end{pmatrix}=3\;.$
The notation is adopted from [2], where the superscript $\!\!\,{}^{(1)}$ marks tangent versions of the original function $f$ and of its variables ${\bf x}$ and $y.$
A (first-order) adjoint version
$f_{(1)}:{\mathbb{R}}^{N}\times{\mathbb{R}}\rightarrow{\mathbb{R}}^{N}:{\bf x}_{(1)}=f_{(1)}({\bf x},y_{(1)})$
resulting from adjoint AD applied to a given differentiable implementation of $f$ computes
${\bf x}_{(1)}\equiv y_{(1)}\cdot f^{\prime}({\bf x})$
without prior accumulation of the gradient. Set $y_{(1)}=1$ to obtain the latter by a single evaluation of $f_{(1)},$ for example, ${\bf x}_{(1)}=1\cdot f^{\prime}({\bf x})=1\cdot(1~{}2~{}3~{}4)=(1~{}2~{}3~{}4).$ As in [2], the subscript $\!\!\,{}_{(1)}$ marks adjoint versions of the original function $f$ and of its variables ${\bf x}$ and $y.$
1.2 The Code
Consider the following implementation of $f$ in C++.
template<int N, typename T>
void f(T x[], T &y) {
  y=0;
  for (int i=0;i<N;i+=2) {
    y+=x[i]*x[i+1];
Instantiation with a given size N for the input vector x with elements of a given type T triggers evaluation of f in T-arithmetic. All operators (=,+=,*) need to be defined for type T, which is certainly the case for all built-in arithmetic types, such as float, int, … as well as char.
Tangent AD yields $f^{(1)}$ as follows:
// tangent version of f
template<int N, typename T>
void f_t1(T x[], T x_t1[], T &y_t1) {
  y_t1=0;
  for (int i=0;i<N;i+=2) {
     // product rule
    y_t1+=x_t1[i]*x[i+1]+x[i]*x_t1[i+1];
Given values of the two input vectors x and x_t1 define the scalar output y_t1. The superscript $\!\!\,{}^{(1)}$ in the mathematical notation is represented by _t1 in the C++ source code. Well-known tangent code generation rules [2], for example, the product rule, are applied.
Adjoint AD yields $f_{(1)}$ as follows:
// adjoint version of f
template<int N, typename T>
void f_a1(T x[], T y_a1, T x_a1[]) {
  for (int i=N-2;i>=0;i-=2) {
     // adjoint product rule
    x_a1[i]=y_a1*x[i+1];
    x_a1[i+1]=y_a1*x[i];
Given values of the two inputs x and y_a1 define the scalar output vector x_a1. The subscript $\!\!\,{}_{(1)}$ in the mathematical notation is represented by _a1 in the C++ source code. Well-known adjoint code generation rules are applied; see also [2].
The gradient is printed to std::cout in tangent mode by the following driver routine:
// gradient in tangent mode
template<int N, typename T>
void dfdx_t1(T x[]) {
   // declare tangent variables
  T x_t1[N], y_t1;
   // prepare for Cartesian basis vectors e_i
  for (int i=0;i<N;++i) x_t1[i]=0;
  for (int i=0;i<N;++i) {
     // set tangent of input equal to e_i
    x_t1[i]=1;
     // call tangent of f
    f_t1<N>(x,x_t1,y_t1);
     // print tangent of output
    std::cout << y_t1;
     // prepare for next Cartesian basis vector
    x_t1[i]=0;
All steps are documented by comments in the source code. Following initializations, N evaluations of the tangent of $f$ are performed with tangents of the input ranging over the corresponding Cartesian basis vectors. The gradient
$f^{\prime}({\bf x})=(77~{}101~{}114~{}114~{}121~{}32~{}88~{}109~{}97~{}115)$
is computed and printed entry by entry.
The same task is completed more efficiently in adjoint mode by the following driver:
// gradient in adjoint mode
template<int N, typename T>
void dfdx_a1(T x[]) {
   // declare adjoint variables
  T x_a1[N]; T y_a1;
   // set adjoint of output equal to one
  y_a1=1;
   // call adjoint of f
  f_a1<N>(x,y_a1,x_a1);
   // print adjoint of input
  for (int i=0;i<N;++i) std::cout << x_a1[i];
A single evaluation of the adjoint function is performed with the adjoint of the output set equal to one. All entries of the gradient are computed simultaneoulsy.
All of the above is stored in the C++ header file f1.h. It is included into the following source of the entire program.
#include ”f1.h”
#include <iostream>
#include <chrono>
int main() {
  using namespace std;
  using namespace std::chrono;
   // increase to experience superior run time of adjoint
  const int N=10;  // required to be even
   // code to be executed in integer arithmetic
  using T=char;
   // size of x equal to size of gradient
  T x[N];
   // vanishing inputs up to entry N-10
  for (int i=0;i<N-10;++i) x[i]=0;
   // custom input values x[N-10]  x[N-1]
  T xv[]={101,77,114,114,32,121,109,88,115,97};
  for (int i=N-10;i<N;++i) x[i]=xv[i-(N-10)];
   // Season’s Greetings
  cout << ”ADwishes;
   // start time measurement
  auto begin=system_clock::now();
   // compute gradient in tangent
  dfdx_t1<N>(x);
   //  or adjoint mode
   // dfdx_a1<N>(x);
   // finish time measurement
  auto end=system_clock::now();
   // report run time
  cout << ”!(taking
    << duration_cast<seconds>(end-begin).count()
    << ”s)” << endl;
  return 0;
The last ten elements of an otherwise zero vector ${\bf x}\in{\mathbb{R}}^{N}$ are set equal to
$(101~{}77~{}114~{}114~{}32~{}121~{}109~{}88~{}115~{}97)^{T}$
and the elapsed run time of the computation of the gradient in tangent or adjoint modes is measured for given even $N\geq 10.$
1.3 The Run Times
The code is compiled using g++ -O3 on my Linux laptop. Both alternatives wish Merry Xmas! Leading zeros in ${\bf x}$ yield zero entries in the gradient. Their output to std::cout has no visible effect due to interpretation as empty 0-terminated C-strings.
Tangent mode takes six seconds to complete the task for $N=231224$. Adjoint mode gets the job done without noticeable delay. You are encouraged to let your own experiments illustrate the $\mathcal{O}(N)$ growth of this gap in run time.
2 AD wishes Happy 2024!
2.1 The Maths
The Hessian of the function $f:{\mathbb{R}}^{N}\rightarrow{\mathbb{R}}:y=f({\bf x}),$
$y=\frac{1}{6}\cdot\sum_{i=0}^{N-1}x^{3}_{i}\;,$
where ${\bf x}=(x_{j})_{j=0}^{N-1}$ and $N\geq 0,$ is easily found to be equal to
$f^{\prime\prime}({\bf x})=\left(h_{j,i}\right)^{i=0,\ldots,n-1}_{j=0,\ldots,n-1},\;\;\text{where}\;\;h_{j,i}=\begin{cases}x_{i}&i=j\\ 0&i\neq j\end{cases}\;.$
For example,
$f^{\prime\prime}({\bf x})=\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&4\end{pmatrix}$
at ${\bf x}=(1~{}2~{}3~{}4)^{T}.$
A second-order tangent version
$f^{(1,2)}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\rightarrow{\mathbb{R}}:y^{(1,2)}=f^{(1,2)}({\bf x},{\bf x}^{(1)},{\bf x}^{(2)})$
of $f$ computes
$y^{(1,2)}\equiv{\bf x}^{(1)^{T}}\cdot f^{\prime\prime}({\bf x})\cdot{\bf x}^{(2)}$
without explicit accumulation of the Hessian. As in [2], we denote tangents due to the application of AD to $f$ by the superscript $\!\!\,{}^{(1)}.$ Reapplication of tangent AD to $f^{(1)}$ appends the superscript $\!\!\,{}^{(2)}$ to the respective variable and function names. Sequences of superscripts are fused into $\!\!\,{}^{(1,2)}\equiv\!\!\,^{(1)^{(2)}}.$
A dense Hessian requires ${\bf x}^{(1)}$ and ${\bf x}^{(2)}$ to range independently over the Cartesian basis vectors in ${\mathbb{R}}^{N}.$ Potential symmetry should be exploited. $f^{(1,2)}$ needs to be evaluated $\frac{N(N+1)}{2}$ times in this case. In the given special case the diagonal structure of $f^{\prime\prime}$ allows for its accumulation with only $N$ evaluations of $f^{(1,2)}$ by letting ${\bf x}^{(1)}$ and ${\bf x}^{(2)}$ range simultaneously over the Cartesian basis vectors in ${\mathbb{R}}^{N}.$ For example,
${\bf e}_{2}^{T}\cdot f^{\prime\prime}({\bf x})\cdot{\bf e}_{2}=\begin{pmatrix}0&0&1&0\end{pmatrix}\cdot\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&4\end{pmatrix}\cdot\begin{pmatrix}0\\ 0\\ 1\\ 0\end{pmatrix}=3$
at ${\bf x}=(1~{}2~{}3~{}4)^{T}.$
A second-order adjoint version
$f_{(1)}^{(2)}:{\mathbb{R}}^{N}\times{\mathbb{R}}\times{\mathbb{R}}^{N}\rightarrow{\mathbb{R}}^{N}:{\bf x}_{(1)}^{(2)}=f_{(1)}^{(2)}({\bf x},y_{(1)},{\bf x}^{(2)})$
of $f$ computes
${\bf x}_{(1)}^{(2)}\equiv y_{(1)}\cdot f^{\prime\prime}({\bf x})\cdot{\bf x}^{(2)}$
without prior accumulation of the Hessian. As in [2], we denote adjoints due to the application of AD to $f$ by the subscript $\!\!\,{}_{(1)}.$ Subsequent application of tangent AD to $f_{(1)}$ appends the superscript $\!\!\,{}^{(2)}$ to the respective variable and function names.
Set $y_{(1)}=1$ and let ${\bf x}^{(2)}$ range over the Cartesian basis vectors in ${\mathbb{R}}^{N}$ to accumulate the Hessian column-wise by $N$ evaluations of $f_{(1)}^{(2)}.$ For example,
$1\cdot f^{\prime\prime}({\bf x})\cdot{\bf e}_{2}=1\cdot\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&4\end{pmatrix}\cdot\begin{pmatrix}0\\ 0\\ 1\\ 0\end{pmatrix}=\begin{pmatrix}0\\ 0\\ 3\\ 0\end{pmatrix}$
at ${\bf x}=(1~{}2~{}3~{}4)^{T}.$
Direct compression [3] of a diagonal Hessian exploits the fact that all columns are structurally orthogonal as no two of them contain nonzero entries in the same row. The sum of all columns amounts to the diagonal. It is obtained by a single evaluation of $f_{(1)}^{(2)}$ with all entries of ${\bf x}^{(2)}$ set equal to one. For example,
$1\cdot f^{\prime\prime}({\bf x})\cdot\sum_{i=0}^{N}{\bf e}_{i}=1\cdot\begin{pmatrix}1&0&0&0\\ 0&2&0&0\\ 0&0&3&0\\ 0&0&0&4\end{pmatrix}\cdot\begin{pmatrix}1\\ 1\\ 1\\ 1\end{pmatrix}=\begin{pmatrix}1\\ 2\\ 3\\ 4\end{pmatrix}$
at ${\bf x}=(1~{}2~{}3~{}4)^{T}.$
2.2 The Code
Consider the following implementation of $f$ in C++.
template<int N, typename T>
void f(T x[], T &y) {
  y=0;
  for (int i=0;i<N;++i) {
    y+=pow(x[i],3);
  y/=6;
Tangent AD yields $f^{(1)}$ as follows:
// first-order tangent version of f
template<int N, typename T>
void f_t1(T x[], T x_t1[], T &y_t1) {
   // initially vanishing derivative of result
  y_t1=0;
  for (int i=0;i<N;++i) {
     // tangent of cubic function; factor 3 pulled out
    y_t1+=pow(x[i],2)*x_t1[i];
   // factor 3 applied
  y_t1/=2;
Analogously, adjoint AD yields $f_{(1)}$ as follows:
// first-order adjoint version of f
template<int N, typename T>
void f_a1(T x[], T y_a1, T x_a1[]) {
  for (int i=N-1;i>=0;–i) {
     // adjoint of cubic function; factor 1/6 pulled in
    x_a1[i]=pow(x[i],2)/2*y_a1;
Application of tangent AD to $f^{(1)}$ yields the second-order tangent $f^{(1,2)}$ as follows:
// second-order tangent version of f
template<int N, typename T>
void f_t1_t2(T x[], T x_t1[], T x_t2[], T &y_t1_t2) {
   // initially vanishing derivative of result
  y_t1_t2=0;
  for (int i=0;i<N;++i) {
     // tangent of quadratic function; factor 1/2 applied
    y_t1_t2+=x[i]*x_t1[i]*x_t2[i];
Analogously, application of tangent AD to $f_{(1)}$ yields the second-order adjoint $f_{(1)}^{(2)}$ as follows:
// second-order adjoint version of f
template<int N, typename T>
void f_a1_t2(T x[], T y_a1, T x_t2[], T x_a1_t2[]) {
  for (int i=N-1;i>=0;–i) {
     // adjoint of quadratic function; factor 1/2 applied
    x_a1_t2[i]=x[i]*y_a1*x_t2[i];
Without exploitation of sparsity the second-order tangent function is called $\frac{N(N+1)}{2}$ times to compute the lower triangular submatrix of the Hessian
$f^{\prime\prime}({\bf x})=\begin{pmatrix}72&0&0&0&0&0&0&0&0&0\\ 0&97&0&0&0&0&0&0&0&0\\ 0&0&112&0&0&0&0&0&0&0\\ 0&0&0&112&0&0&0&0&0&0\\ 0&0&0&0&121&0&0&0&0&0\\ 0&0&0&0&0&32&0&0&0&0\\ 0&0&0&0&0&0&50&0&0&0\\ 0&0&0&0&0&0&0&48&0&0\\ 0&0&0&0&0&0&0&0&50&0\\ 0&0&0&0&0&0&0&0&0&52\\ \end{pmatrix}\;.$
The following driver computes the individual entries of the diagonal by letting ${\bf x}^{(1)}$ and ${\bf x}^{(2)}$ range simultaneously over the Cartesian basis vectors in ${\mathbb{R}}^{N}.$
// Diagonal Hessian in second-order tangent mode
template<int N, typename T>
void ddfdxx_t1_t2_sparse(T x[]) {
   // declare tangent variables
  T x_t1[N], x_t2[N], y_t1_t2;
   // prepare for Cartesian basis vectors e_i
  for (int i=0;i<N;++i) x_t1[i]=x_t2[i]=0;
   // loop over diagonal entries
  for (int i=0;i<N;++i) {
     // set both tangents of inputs equal to e_i
    x_t1[i]=x_t2[i]=1;
     // call second-order tangent version of f
    f_t1_t2<N>(x,x_t1,x_t2,y_t1_t2);
     // print i-th diagonal entry of Hessian
    std::cout << y_t1_t2;
     // prepare for next Cartesian basis vector
    x_t1[i]=x_t2[i]=0;
The second-order tangent function is called $N$ times.
Direct compression [3] in second-order adjoint mode yields the sum of all columns of the Hessian as follows:
// Column-compressed Hessian in second-order adjoint mode
template<int N, typename T>
void ddfdxx_a1_t2_compressed(T x[]) {
   // declare tangent and adjoint variables
  T x_t2[N], x_a1_t2[N], y_a1;
   // set all entries of tangent of input equal to one
  for (int i=0;i<N;++i) x_t2[i]=1;
   // set adjoint of output equal to one
  y_a1=1;
   // call second-order adjoint f
  f_a1_t2<N>(x,y_a1,x_t2,x_a1_t2);
   // print sum of all columns of Hessian
  for (int j=0;j<N;++j) std::cout << x_a1_t2[j];
A single call of the second-order adjoint function yields all diagonal entries of the Hessian.
All of the above is stored in the C++ header file f2.h. It is included into the following source of the entire program.
#include ”f2.h”
#include <iostream>
#include <chrono>
int main() {
  using namespace std;
  using namespace std::chrono;
   // increase to compare run times of various approaches
  const int N=240101;
   // code to be executed in integer arithmetic
  using T=char;
   // size of x yields squared number of Hessian entries
  T x[N];
   // vanishing inputs up to entry N-10
  for (int i=0;i<N-10;++i) x[i]=0;
   // custom input values x[N-10]  x[N-1]
  T xv[]={72,97,112,112,121,32,50,48,50,52};
  for (int i=N-10;i<N;++i) x[i]=xv[i-(N-10)];
   // Season’s Greetings
  cout << ”ADwishes;
   // start time measurement
  auto begin=system_clock::now();
   // compute diagonal of the Hessian in tangent
  ddfdxx_t1_t2_sparse<N>(x);
   //  or adjoint mode
   // ddfdxx_a1_t2_compressed<N>(x);
   // finish time measurement
  auto end=system_clock::now();
   // report run time
  cout << ”!(taking
    << duration_cast<seconds>(end-begin).count()
    << ”s)” << endl;
  return 0;
The last ten elements of an otherwise zero vector ${\bf x}\in{\mathbb{R}}^{N}$ are set equal to
$(72~{}97~{}112~{}112~{}121~{}32~{}50~{}48~{}50~{}52)^{T}$
and the elapsed run time of the computation of the diagonal of the Hessian in sparsity-aware second-order tangent or adjoint modes is measured for given $N\geq 10.$
2.3 The Run Times
Both alternatives wish Happy 2024!
On my laptop, sparse second-order tangent mode takes nine seconds to complete this task for $N=240101$. Direct compression in second-order adjoint mode gets the job done almost immediately. Again, you are encouraged to let your own experiments illustrate the $\mathcal{O}(N)$ growth of this gap in run time.
3 Conclusion
Merry Xmas and Happy 2024!